OpenStudy (anonymous):

When a product is raised to a power, it is equal to the product of each factor raised to that power. Show that this rule does not apply for a sum or difference to a power using the problem (5 – 3)2

OpenStudy (anonymous):

5^2-3^2=? (5-3)^2=2^2=? Are they the same?

OpenStudy (anonymous):